Opened 7 years ago

Closed 5 years ago

Last modified 16 months ago

#441 closed flight processing (fixed)

GB12/06, flight day 172a/2012, Harnhill

Reported by: knpa Owned by: knpa
Priority: alpha 4 medium Milestone: 2012 data processing completion
Component: Processing: general Keywords:
Cc: Other processors: adbe

Description (last modified by adbe)

Data location: /users/rsg/arsf/arsf_data/2012/flight_data/uk/GB12_06-2012_172a_Harnhill

Data arrived from ARSF via network transfer on 22/06/2012

GB12/06 Scientific objective:

GB12/06 Priority: a4m

PI: Anthony Beck

Sensors:

Camera (6/7/2012)
Eagle (6/7/2012)
Hawk (6/7/2012)
lidar (+FWF) (6/7/2012)

Change History (27)

comment:1 Changed 7 years ago by knpa

  • Description modified (diff)
  • Summary changed from GB12/06, flight day 201a/2012, Harnhill to GB12/06, flight day 172a/2012, Harnhill

comment:2 Changed 7 years ago by knpa

  • Description modified (diff)

comment:3 Changed 7 years ago by knpa

  • Description modified (diff)
  • Owner set to knpa
  • Status changed from new to assigned

This flight was incorrectly labelled 201 by ops, so it still says this on the handwritten logsheet and possibly in other places.

comment:4 Changed 7 years ago by adbe

Processed navigation. When processing in grafnav there was a drop out of the navigation on a bend around the flight area, Mark thinks that this could be due to losing satellites when banking; hopefully this will not cause any problems later on.

comment:5 Changed 7 years ago by adbe

Started processing hyperspectral.

comment:6 Changed 7 years ago by adbe

Found sct's

line eagle hawk
1 -0.04 -0.02
2 -0.04 -0.02
3 -0.04 -0.06
4 -0.04 -0.04
5 -0.04 -0.04
6 -0.04 -0.08
7 -0.04 -0.06

comment:7 Changed 7 years ago by adbe

Hyperspectral delivery created.

comment:8 Changed 7 years ago by knpa

adbe has created the rcd tiffs, thumbnails and processed image events file.

comment:9 Changed 7 years ago by knpa

  • Other processors set to adbe

comment:10 Changed 7 years ago by knpa

Removed tiffs 1-16 as these are calibration ones.

comment:11 Changed 7 years ago by knpa

Camera ready for delivery check.

comment:12 Changed 7 years ago by adbe

Camera delivery check:

  • ARSF Project Code: GB12_06 (should be GB12/06)

Otherwise, ready to deliver.

comment:13 Changed 7 years ago by knpa

Fixed proj code format, ready to deliver.

comment:14 Changed 7 years ago by knpa

Beginning LiDAR processing.

adbe has already created the las files (incl. full waveform)

comment:15 Changed 7 years ago by knpa

No vertical or horizontal discrepancies found.
Lines up well with OS vectors.

comment:16 Changed 7 years ago by knpa

las files fully classified.

comment:17 Changed 7 years ago by knpa

Beginning HS delivery check.

comment:18 Changed 7 years ago by knpa

Delivery check complete. No problems found.

The only thing is strange is that the 3rd line is so short (it doesn't even come up to the cross line) and this gives a gap in coverage. Compare this with the lidar where that lines goes all the way across the site. Maybe look into this and see if there was a problem somewhere, or if the instrument was just switched off at that point.

comment:19 Changed 7 years ago by knpa

adbe has done the delivery for LiDAR so this is ready for delivery check (not by me or adbe)

comment:20 Changed 7 years ago by adbe

The hand-written logsheet mentioned that the Eagle and Hawk had mis-triggered for the line 3. So have added this to typed logsheet and ReadMe.

comment:21 Changed 7 years ago by jaho

LiDAR delivery check:

No issues. Ready to deliver.

comment:22 Changed 7 years ago by adbe

  • Description modified (diff)

Camera, Hyperspectral, Lidar(+FWF) delivered to Anthony Beck on 6/7/2012 on disk 124.

comment:23 Changed 6 years ago by knpa

  • Component changed from Processing: general to Archiving

test

Last edited 6 years ago by knpa (previous) (diff)

comment:24 Changed 6 years ago by knpa

  • Component changed from Archiving to Processing: general

comment:25 Changed 5 years ago by tipo

Prepared for archiving (27/11/2013)

comment:26 Changed 5 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Archived by NEODC.

Closing ticket.

comment:27 Changed 16 months ago by asm

JSON files - Lidar delivery

Problem with bng dem projection missing, solved using assign_projection.py -p '+proj=tmerc +lat_0=49 +lon_0=-2 +k=0.9996012717 +x_0=400000 +y_0=-100000 +ellps=airy +units=m +no_defs ' the_dem-bng.dem

Note: See TracTickets for help on using tickets.